9x+3y= -48 -6x+y= 38
Found 2 solutions by fractalier, reelmccray:
Answer by fractalier(6550)

I would divide the first equation by three and subtract the second equation from it, like this
3x + y = -16
-(-6x + y = 38)
---------------------
9x = -54
x = -6
Then substitute back in and get
3(-6) + y = -16
-18 + y = -16
y = 2

Answer by reelmccray(6)

I assume you want to solve for X and Y.
9X + 3Y = -48
solve for Y
3Y = -9X - 48
Y = -3X - 16
since (-3X - 16) = Y
Substitute (-3X - 16) for Y in -6X + Y = 38
which will look like
-6X + (-3X - 16) = 38
Solve for X
-9X = 54
X = -6
Now that you have determined X = -6 use the -6 for X in one of the equations to find Y.
-6X + Y = 38
-6(-6) + Y = 38
36 + Y = 38
solve for Y
Y = 2